OpenStudy (anonymous):

Evaluate the expressions for x=3 and y=2 -4x+3y

OpenStudy (compassionate):

-4(3) + 3(2) Can you solve it from here?

OpenStudy (anonymous):

okay so -4(3) IS 12. but would it be negative 12.

OpenStudy (anonymous):

-4x+3y x=3 and y=2 -4(3)+3(2) -12 + 6 . . .

OpenStudy (anonymous):

when multiplying: negative * positive = negative

OpenStudy (anonymous):

-4 * 3 = -12